OpenStudy (anonymous):

Write a quadratic equation win the form ax2 +bx + c=0, where a,b,c are integers for the give roots. X=3 or x =-5.

OpenStudy (campbell_st):

if you are given the roots x = 3 then x - 3 = 0 so (x - 3) is a binomial factor and x = -5 then x + 5 = 0 so (x + 5) is a binomial factor multiply the 2 binomial factors for the quadratic

OpenStudy (anonymous):

Thank you sooo sooo much!!